OpenStudy (anonymous):

Solve\[\int_0^\infty\frac{\ln\left(x+1\right)}{x}\]

OpenStudy (anonymous):

Just kidding, I'm really here to mess around with LaTeX.

OpenStudy (anonymous):

\[\huge \int\limits_{\underbrace{0}_{\text{Can't solve it !}}}^{\overbrace{\infty}^{\text{Too Far!}} }=\cfrac{\ln (x+1)}{\cfrac{x}{\cfrac{1}{\cfrac{1}{\cfrac{1}{\cfrac{1\cdot 1^5}{\cfrac{400^0\cdot 1^7}{1}}}}}}}\] LOOOL...

OpenStudy (anonymous):

\[\huge{\text{OH LOOK AT YOU, SO SPECIAL.}}\]

OpenStudy (anonymous):

hahaah... LOL

OpenStudy (anonymous):

\[\large{\text{I KNOW MY /UNDERBRACE SMARTYPANTS}}\]

OpenStudy (anonymous):